OpenStudy (anonymous):

The measure of the length of a field is 7x + 2 yards and the measure of the width of the field is 4 + x yards. Write an expression for its perimeter. 8x + 6 14x 16x + 12 28x

OpenStudy (lalaly):

perimeter = 2(length+width)

OpenStudy (anonymous):

P = 2(length + width) P = 2(7x+2 + 4 + x) P = 2(8x + 6) P = 16x + 12 Final answer: 16x + 12 yards